Vogue Merchants uses a periodic inventory system. The beginning inventory of a particular product and the purchases during the current year, were as follows Jan 1 Mar9 Purchase Aug 11 Purchse Dec 23Purchasre Begnning nventory 1,000 units $13.00$13.0o 1,050 units $13 50 14,175 950 unats $14.00 13,300 Total avalable for sale n year 500 unats $14 75 500 units At December 31, the ending inventory of this product consisted of 850 units. Determine the cost of the year-end inventory and the cost of goods sold for this product under each of the following methods of inventory valuation (Rounded): Inventory atCost of Dec. 31 Goods S Average cost First-in, first-out Last $1 n, first-out
Answer #1

Note : Unit sold = 3,500 - 850 = 2,650 units

Answer (a)

Value of ending inventory under Average cost method = 850 units * ($47,850 / 3,500 units) = $11,620

Value of cost of goods sold under Average cost method = $47,850 - $11,620  = $36,230

Answer (b)

Value of ending inventory under First in , first - out method

= (350 * $14) + (500 * $14.75) = $12,275

Value of cost of goods sold under First in , first - out method

= (1,000 * $ 13) + (1,050 * $13.50) + (600 * $14) = $35,575

Answer (c)

Value of ending inventory under Last in , first - out method

= (850 * $13) = $11,050

Value of cost of goods sold under Last in , first - out method

= (500 * $14.75) + (950 * $14) + (1,050 * $13.50) + (150 * $13)= $36,800

Summary :

Inventory at Dce 31 Cost of goods sold at Dec 31
Average cost $11,620 $36,230
First in , first - out $12,275 $35,575
Last in , first - out method $11,050 $36,800